The temperature in Frostport is 8 degrees & dropping @ the rate of 3.5 degrees per hour. How many hours will it take for the temperature to reach -20 degrees.
Answer by ikleyn(52792)

.
The distance between these two marks in number line is  8 - (-20) = 8 + 20 = 28.


To answer the question, divide 28 by 3.5.

That's all.